OpenStudy (anonymous):

A motel bills its customers by charging a flat fee for the first night and then adding on a fixed amount for every night thereafter. If it costs George $155 to stay in the motel for 3 nights and Noah $290 to stay in the motel for 6 nights, how much is the flat fee for the first night? A motel bills its customers by charging a flat fee for the first night and then adding on a fixed amount for every night thereafter. If it costs George $155 to stay in the motel for 3 nights and Noah $290 to stay in the motel for 6 nights, how much is the flat fee for the first night? @Mathematics

OpenStudy (anonymous):

x+2y=155 x+5y=290 solve system of equations and x will be your one night payment

OpenStudy (anonymous):

It's a linear function y = mx + b

OpenStudy (anonymous):

above are linear equations?!

OpenStudy (anonymous):

you have two fixed amounts so you don't have to worry about exponential growths or anything

OpenStudy (anonymous):

yeah so 3y =135 y = 45 x = 65

OpenStudy (anonymous):

so the flat fee is 65

OpenStudy (anonymous):

x is the flat fee and y is the rate or slope

OpenStudy (anonymous):

if you want to think of it that way i guess lol